The electric potential at  any point x,y, and z in metres is given by V=3x2 .the electric field at a point (2,0,1)   is

Grade:12

1 Answers

Arun
5 years ago
 
Electric potential V = 3x2 volts
The electric potential changes only along x-axis, We know that Ex = -dV/dx? Ex = - d(3x2)/ dx = - 6x
The electric field at point (2, 0, 1) will be (here x = 2) Ex  = – 12 volt/m
 
 
Regards
Arun (askIITians forum expert)